OpenStudy (anonymous):

please let f(x)= x [ 1/2 (1-(X^2))^1/2 (-2x) ] + sqroot (1-(x^2)) a. For what values is f(x)=0? b. For what values is f(x) undefined? c. For what values is f(x)<0? d. For what values is f(x)>0?

OpenStudy (anonymous):

\[f(x) = x[\frac{1}{2}(1-x^2)^{\frac{1}{2}}(-2x)]+\sqrt{1-x^2}\] is this the equation?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[\text{A. }\ x = 1\ \ or\ \ x = -1\]

OpenStudy (anonymous):

\[x \left[ \frac{1}{2} (1-x^2)^\frac{1}{2} (-2x) \right] + \sqrt{1-x^2} \]\[ = x \left[ - x\sqrt{(1-x^2)} \right] + \sqrt{1-x^2} \]\[ = -x^2\sqrt{1-x^2} + \sqrt{1 - x^2}\]\[ = (1 - x^2)\sqrt{1 - x^2}\]\[ = (1-x^2)^{\frac{3}{2}}\] Part A: \[(1-x^2)^{\frac{3}{2}} = 0\]\[x^2 = 1\]\[x = \pm1\] Part B: f(x) is undefined when: \[(1 - x^2) < 0\]\[x^2 > 1\]\[x < -1 \space\vee\space x > 1\]

OpenStudy (anonymous):

\[\text{B. }\ x >1\ \ or\ \ x < 1\] \[\text{C. }\phi\] \[\text{D. } x < 1\ or\ x > -1\]

OpenStudy (anonymous):

in other words for D. -1 < x < 1

OpenStudy (anonymous):

Part C: I think no values of x satisfy this, as the square root will only give positive values: Part D: Any values between -1 and 1 exclusive. Because 1 - x^2 must be greater than 0.

OpenStudy (anonymous):

thanks a lot . appreciate it